An automobile accelerates from rest at 1.90 m/s2 for 19.6 s. The speed is then held constant for 19.6 s, after which there is an acceleraton of -2.82 m/s2 until the automobile stops. What is the total distance traveled? Show work.
Answer #1

Using the kinematic equation: S = ut + . = 0+ (1.90)(19.6) = 364.95 m U2 = uz + at = 0+(1.9) (19.6) = 37.24 m/s S, = ut = (37